OpenStudy (anonymous):

ALGEBRA 2 HELP EASY! (I GIVE MEDALS!) What is the solution set in interval notation for |3x + 1| <= 7

OpenStudy (anonymous):

Your an idiot this is common sense

OpenStudy (lxelle):

3x+1 =-7 and 3x+1=7 solve it from there.

OpenStudy (anonymous):

I got (-infinity, infinity) as the answer, but it was marked wrong.

OpenStudy (anonymous):

@LXelle

OpenStudy (anonymous):

Do you guys like tacos

OpenStudy (lxelle):

How did you get infinity? Loll.

OpenStudy (anonymous):

WHY ARE YOU IGNORING ME

OpenStudy (anonymous):

for the first one I got x <= 2 and x >= -8/3

OpenStudy (anonymous):

Sorry for calling you an idiot

OpenStudy (anonymous):

and when you graph that it takes up the entire line

OpenStudy (anonymous):

Elmo's sorry

OpenStudy (anonymous):

@LXelle

OpenStudy (anonymous):

I hate acacia Brinkley tbh

OpenStudy (anonymous):

same lmao

OpenStudy (anonymous):

dont feel like changing my pic tho

OpenStudy (anonymous):

I'm in class rn pretending to do math

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

Some kids beat boxing across the room

OpenStudy (anonymous):

Can you help me with a math problem by chance

OpenStudy (anonymous):

Lol sure

OpenStudy (lxelle):

No, we are not finding the domain. We are just finding the range of x.

OpenStudy (lxelle):

What you got there is correct.

OpenStudy (anonymous):

What would the solution set in interval notation be for -|3x + 5| < 4?

OpenStudy (lxelle):

Do like what you did just now.

OpenStudy (anonymous):

so I got x > 1/-3 and x < -1/3 is that right?

OpenStudy (anonymous):

So how would we write that in interval notation?

OpenStudy (lxelle):

x= -3 and x=-1/3 check your answers up there.

OpenStudy (anonymous):

Right sorry!

OpenStudy (anonymous):

So would we write that as [-3,-1.3]?

OpenStudy (anonymous):

*-1/3

OpenStudy (lxelle):

It would be x>-13 and x>-3

OpenStudy (anonymous):

-13?

OpenStudy (lxelle):

SORRYY I MEAN -1/3

OpenStudy (anonymous):

Oh lol it's fine

OpenStudy (anonymous):

Wouldn't it be x < -1/3? or do you flip the sign when dividing a negative number?

OpenStudy (lxelle):

-3x-5<-4, -3x<-4+5, -3x<1, x>-1/3

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

But how would we write the final answer in interval notation?

OpenStudy (anonymous):

@LXelle

OpenStudy (lxelle):

Not sure with that. Might not want to confuse you. Sorry.

OpenStudy (anonymous):

Okay it's fine! Thanks for your help though!

OpenStudy (lxelle):

:))